# Algebraic geometry of bubbling Kähler metrics

### v1

@@ Abstract.
We give an algebro-geometric or non-archimedean framework to study bubbling phenomena of Kähler metrics with Euclidean volume growth, after [DS17, Sun23, dBS23]. In particular, for any degenerating family to log terminal singularity, we algebraically construct a finite sequence of birational modifications of the family with milder degenerations, and compare with analytic bubbling constructions in [Sun23, dBS23]. We also provide approaches in terms of coordinates and valuations. Our discussion partially depends on the general framework of stability theory in [Od24b] after [HL14, AHLH23].
@@ 1. Introduction
We make a purely algebro-geometric study of degenerations of normal singularities in this paper, motivated by the recent works on asymptotic behaviour of Kähler metrics [DS17, Sun23, dBS23]. Our main setup is as follows.
@@
We also assume $K_{\mathcal{X}}$ to be $\mathbb{Q}$-Cartier as our main interest comes from Kähler-Einstein geometry. We set $X:=\mathcal{X}_{0}:=\pi^{-1}(0)$. We further fix a holomorphic (resp., algebraic) section $\sigma$ with $\sigma(0)=x$.
In the next section, we algebraically construct (almost) canonical, finite ordered sequences of modifications of $\mathcal{X}$ with milder degenerations. For instance, if the general fibers are $2$-dimensional smooth, degenerating to $X$ with a $2$-dimensional $A_{k}$-singularity, then we obtain other fillings with finite number of $A_{i}$-singularities with smaller Milnor number $i$ (cf., [dBS23]) and terminate at smooth filling. The process is different from the classical simultaneous resolution of Brieskorn-Grothendieck type in an essential manner. Morally speaking, our algorithm gives canonical “limit singularities” of various rescales beyond the given $X$ itself or its (metric) tangent cone, reflecting the whole formation process of the singularities of $X$, in the context of K-stability. While we explain its detailed meaning later, here is the brief overview of the differential geometric background.
Suppose we have a “good” family of (weak) Kähler metrics $g_{t}$ on $\mathcal{X}_{t}=\pi^{-1}(t)$ in the pluri-potential theoritic sense (cf., [BBEGZ19]), for $t\neq 0$, especially of Kähler-Einstein metrics, we would like to consider and answer the following question:
Main Question (see [Sun23, dBS23]).
@@
In this paper, following [Sun23, dBS23], we call such a non-collapsing pointed Gromov-Hausdorff limit simply as bubbles of the family $(\mathcal{X}_{t},g_{t})$ for $t\to 0$. By the groundbreaking work [DS17, 1.4], extending some classical cases, such limits often naturally admits structure of an affine log terminal variety with (weak) Ricci-flat Kähler metric. Indeed, our aforementioned main construction of affine varieties is expected to underlie these bubbles. They have Euclidean volume growth simply by the Bishop’s inequality. This is a (Kähler) metric version of so-called bubbling phenomenon since [SU81], as later also studied by [Nak88, And89, Ban90, BKN89] in mainly $4$-dimensional Riemannian geometric setup, which in particular found out the finite energy (the $L^{2}$-norm of the curvature) rescaling provides smooth asymptotically locally Euclidean (ALE) manifolds limits as the deepest bubbles. Then, recently we witnessed a breakthrough [DS17, Sun23, dBS23] in the higher dimensional Kähler setup. In particular, [Sun23] proved the existence of “minimal” bubble in a differential geometric sense, finiteness of bubbles, and raised Conjecture 5.4 in loc.cit to expect a $2$-step algebraic construction of the underlying varieties of the minimal bubbles, and [dBS23] showed many examples and a refinement of the conjecture.
@@
The main contents of this paper give answer the above main question by solving large parts of the conjectures of [Sun23, dBS23] with some refinements. The following is a brief outlook of our statements but see the detailed statements in later sections.
@@
    (Theorem 2.4) There is a purely algebro-geometric $2$-step construction 222the following symbol $\rightsquigarrow$ means each step of the construction, and is not a map. of algebro-geometric version of minimum bubbles
@@
    Here, the minimality morally means the mildness of singularity measured by the normalized volumes (at the base points) as we clarify later, as it is parallel to minimality of the diverging order of the rescaling constants $\{c_{t}\}_{t}$ (cf., also [Sun23] and Definition 2.1).
@@
    (Theorem 2.4, Theorem 2.8) The first step construction
@@
    is uniquely (canonically) characterized, and coincides with more analytic construction following [Sun23, §4], which is called algebraic minimal bubble in [Sun23, §5.1, Conjecture 5.4].
  3. ($\mathrm{iii}$)
    (Theorem 2.6) Even at the total space level, the $1$-st step process $\mathcal{X}_{t}(t\to 0)\rightsquigarrow\mathcal{X}^{\prime}_{\rm min,0}$ can be achieved as a canonical birational modification of the original family, after a possibly non-finite type 333as far as we wants canonicity of the process base change i.e., a certain Novikov type ring spectrum (or an affine toric variety or even $\Delta$, if we do not mind non-canonicity of the base). We refer to these obtained total spaces as $\mathcal{X}^{\prime}_{\rm min}$ and $\mathcal{X}^{\prime\prime}_{\rm min}$.
@@
To make sense of and prove the above ($\mathrm{iii}$), we use [Od24b] in which some algebraic (partly stack-theoretic) generalities to treat degenerations in “irrational direction” with higher dimensional algebraic base is prepared. We also compare our construction with more analytic and differential geometric constructions in [Sun23].
Since our construction is fairly general and look abstract first, although it is done after various concrete calculations as experiments, let us review a simple example after [dBS23] here. In this case, we only need a single process rather than the $2$-step processes i.e., $\mathcal{X}^{\prime}_{\rm min}=\mathcal{X}^{\prime\prime}_{\rm min}$, and they are achieved only after a finite base change of $\Delta$.
Example 1.2 (cf., [BKN89, Ban90] and [dBS23, §3.3]).
@@
We can consider a family of hyperKähler metrics on each $t$ by the Gibbons-Hawking ansatz with the base $\mathbb{R}^{3}$ and monopole points $x_{1},x_{2},x_{3},x_{4}$ and harmonic functions of the form $\frac{1}{2}\sum_{i}\frac{1}{|x-x_{i}|}$. In this case, [dBS23, §3.3] describes how it bubbles at $t\to 0$ around the origin, and coincides with the algebraic procedure Theorem 1.1, and is explicit as follows. (Indeed, our general construction is originally obtained as a generalization of their examples).
@@
after replacing $\frac{x}{t^{2}}$ as $x$, $\frac{y}{t^{2}}$ as $y$, $\frac{z}{t}$ as $z$ while keeping $t$. This is a degeneration to $xy=z^{3}(z-1)$ when $t\to 0$, with $A_{2}$-singularity at $(0,0,0,0)\in\Delta_{t}\times\mathbb{C}_{x,y,z}^{3}$. This admits a ALE hyperKähler orbifold metric, again obtained via the Gibbons-Hawking ansatz over $\mathbb{R}^{3}$ (cf., [dBS23, §3.3]).
@@
We make this kind of explicit study as Example 1.2 via coordinates more generally. That is, after the proof of Theorem 1.1 ($=$Theorems 2.4, 2.8), we propose more explicit ways of constructions of the candidates of bubbles, via coordinates/equations and valuations respectively in §2.4, §2.5. We conjecture that they coincide with the construction of Theorem 2.4. §2.2.3 and §2.3 may be of own independent interest, which are respectively the systematic treatment of negative weight deformation originated in [CH22, SJZ23] and an analogue of bubbling Calabi-Yau metric in the context of Kähler-Ricci flow. (The latter gives a yet another approach to the properness of K-moduli of Fano varieties cf., Theorem 2.21). Further, in §3, we discuss more on the structure on the “bubble tree” formed by all possible bubbles.
@@ 2. Minimal bubble
2.1. Definitions and Review
We work in Setup 1. The main objective of this section is to give an algebraic approach to the “minimum bubble” in the sense of [Sun23, dBS23], whereas the minimality roughly refers to that of the order of rescaling parameters $c_{t}$ for $t\to 0$. We review the definition for convenience.
@@ 2.1. Definitions and Review
In the context of Setup 1, its minimal bubble refers to the pointed Gromov-Hausdorff limits of $(\sigma(t)\in\mathcal{X}_{t},c_{t}^{2}g_{t})$ for certain diverging sequence $c_{t}>0$, which is not a Riemannian metric cone, but whose tangent cone at infinity ([DS17, §3.4]) is the metric local tangent cone at $\sigma(0)\in X$, which we denote as $C_{\sigma(0)}(X)$.
Firstly, we discuss several a priori different algebraic approaches to it constructions, in different subsections, and will discuss the conjectural equivalence later. Since we crucially use the theory of local normalized volume of log terminal singularities due to [Li19, Liu18] (cf., also [Od15b]), with its origin in [DS17], we briefly review the definition and properties we use here. Its smooth case can be traced back to much earlier [Mus02, dFEM04] (and also related [ELS03]) and more detailed review are available at [LLX20], [Od24a, §2] for instance.
@@
Then, the normalized volume $\widehat{{\rm vol}}_{x}(v)$ is defined as $A_{X}(v)^{n}{\rm vol}_{x}(v)$ by [Li19], where $A_{X}(v)$ is the log discrepancy in the generalized sense by [JM12]. Then the following is known by now. We do not make any substantial contribution in this paper but rather use in this paper. See more references in [LLX20], [Od24a, §2].
Theorem 2.3 (cf., [DS17, Li19, Liu18, Blu18, HS17, LX18, LWX21, LLX20, XZ21, Od24a]).
For above $x\in X$, there is a unique quasi-monomial valuation $v=v_{X}$ up to the $\mathbb{R}_{>0}$-action, which minimizes the normalized volume $\widehat{\operatorname{vol}}(-)$.
@@
    Then, the valuation $v_{X}$ coincides with the vanishing order function in [DS17, §3.2] for the local Kähler metric in the setup of loc.cit. We denote the rational rank of $v_{X}$ as $r$. Further, $\widehat{\rm vol}(v_{X})$ has a differential geometric meaning as the volume density of the link of the local metric tangent cone. We denote the minimized value $\widehat{\rm vol}(v_{X})$ as $\widehat{\rm vol}(x\in X)$ or $\widehat{\rm vol}_{X}(x)$, in this paper.
@@
    From more algebraic viewpoint, the graded ring ${\rm gr}_{v}(\mathcal{O}_{X,x}):=\oplus_{s\in{\rm Im}(v_{X})}(\{f\in\mathcal{O}_{X,x}\mid v(f)\geq s\}/\{f\in\mathcal{O}_{X,x}\mid v(f)>s\})$ is a finite type $\mathbbm{k}$-algebra on which there is a natural action of the algebraic $\mathbbm{k}$-torus $T\simeq\mathbb{G}_{m}^{r}$ on a K-semistable Fano cone $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v}(\mathcal{O}_{X,x})$.
@@
We note that the last $\mathbb{Q}$-Gorensteinness assertion of the positive weight degeneration in ($\mathrm{iii}$) seems missing in literature but can be confirmed as follows. As in [Od24b, §2], take homogeneous generating system $\overline{z}_{1},\cdots,\overline{z}_{l}$ of ${\rm gr}_{v}(\mathcal{O}_{X,x})$ with degrees $w_{1},\cdots,w_{l}$, and lift to $\Gamma(\mathcal{O}_{X})$ as $\overline{z}_{1},\cdots,\overline{z}_{l}$ which we suppose to give an embedding $W\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}$ (otherwise add more $\overline{z}_{i}$s). Then we take rational approximation $\xi^{\prime}$ of $\xi:=(w_{1},\cdots,w_{l})$ and corresponding positive weight test configuration as $W_{\xi^{\prime}}$ over $\mathbb{A}^{1}$. For any $\xi^{\prime}$ close enough to $\xi$, we see that the central fiber is again $W$ just with a perturbed vector field induced on it. Note that there is a natural relative compactification $W_{\xi^{\prime}}\subset\overline{W_{\xi^{\prime}}}\to\mathbb{A}^{1}$ by adding the quotient by $\mathbb{G}_{m}$ for $\xi^{\prime}$-direction, which is automatically a $\mathbb{Q}$-Gorenstein family by the arguments of [OSS16, (proof of) Lemma 2.4]. Since this holds for any $\xi^{\prime}$ close enough to $\xi$, the map from the toric base $U_{\tau}$ of the positive weight degeneration is in the $\mathbb{Q}$-Gorenstein deformation component so that we obtain the assertion.
2.2. Re-construction via moduli-theoretic arguments
In this subsection, we use theorems in [Od24b] after [Lan75, HL14, AHLH23] to algebraic construction of the minimal bubble and discuss it. We also discuss reformulation of negative deformations in the sense of [CH22], after [Od24b].

### v2

@@ Abstract.
We give an algebro-geometric or non-archimedean framework to study bubbling phenomena of Kähler metrics with Euclidean volume growth, after [DS17, Sun23, dBS24]. In particular, for any degenerating family to log terminal singularity, we give an algorithm to algebraically construct a finite sequence of birational modifications of the family with milder degenerations, and partially relate with analytic bubbling constructions in [Sun23, dBS24].
On the way, we also introduce and discuss a global analogue which in particular reduces compactness of the moduli of compact Kähler-Ricci solitons, again to a finite generation type problem in birational geometry. Our whole discussion depends on the general framework of stability theory for irrational type degenerations in [Od24b].
@@ 1. Introduction
We make a purely algebro-geometric study of degenerations of normal singularities in this paper, motivated by the recent works on asymptotic behaviour of Kähler metrics [DS17, Sun23, dBS24]. Our main setup is as follows.
@@
Occasionally, to make the base coordinate $t$ clear, we denote $\Delta$ as $\Delta_{t}$. We also assume $K_{\mathcal{X}}$ to be $\mathbb{Q}$-Cartier as our main interest comes from Kähler-Einstein geometry. We set $X:=\mathcal{X}_{0}:=\pi^{-1}(0)$. We further fix a holomorphic (resp., algebraic) section $\sigma$ with $\sigma(0)=x$.
In the next section, we algebraically construct (almost) canonical, finite ordered sequences of modifications of $\mathcal{X}$ with milder degenerations. For instance, if the general fibers are $2$-dimensional smooth, degenerating to $X$ with a $2$-dimensional $A_{k}$-singularity, then we obtain other fillings with finite number of $A_{i}$-singularities with smaller Milnor number $i$ (cf., [dBS24]) and terminate at smooth filling. The process is, weakly related to but substantially different from the classical simultaneous resolution of Brieskorn-Grothendieck type in an essential manner. Morally speaking, our algorithm gives canonical “limit singularities” of various rescales beyond the given $X$ itself or its (metric) tangent cone, reflecting the whole formation process of the singularities of $X$, in the context of Kähler-Einstein geometry and the K-stability. While we explain its detailed meaning later, here is the brief overview of the differential geometric background.
Suppose we have a “good” continuous family of (singular) Kähler metrics $g_{t}$ on $\mathcal{X}_{t}=\pi^{-1}(t)$ in the pluri-potential theoritic sense (cf., e.g., [BBEGZ19]), for $t\neq 0$, especially of Kähler-Einstein metrics, we would like to consider and answer the following question:
Main Question (see [Sun23, dBS24]).
@@
In this paper, following [Sun23, dBS24], we call such a non-collapsing pointed Gromov-Hausdorff limit simply as bubbles of the family $(\mathcal{X}_{t},g_{t})$ for $t\to 0$. By the groundbreaking work [DS17, 1.4], extending some classical cases, such limits often naturally admits structure of an affine log terminal variety with (weak) Ricci-flat Kähler metric. Indeed, our aforementioned main construction of affine varieties is expected to underlie these bubbles. They have Euclidean volume growth simply by the Bishop’s inequality. This is a (Kähler) metric version of so-called bubbling phenomenon since [SU81], as later also studied by [Nak88, And89, Ban90, BKN89] in mainly $4$-dimensional Riemannian geometric setup, which in particular found out the finite energy (the $L^{2}$-norm of the curvature) rescaling provides smooth asymptotically locally Euclidean (ALE) manifolds limits as the deepest bubbles. Then, recently we witnessed a breakthrough [DS17, Sun23, dBS24] in the higher dimensional Kähler setup. In particular, [Sun23] proved the existence of “minimal” bubble in a differential geometric sense, finiteness of bubbles, and raised Conjecture 5.4 in loc.cit to expect a $2$-step algebraic construction of the underlying varieties of the minimal bubbles, and [dBS24] showed many examples and a refinement of the conjecture.
@@
The main contents of this paper give answer the above main question by solving large parts of the conjectures of [Sun23, dBS24] with some refinements. The following is a brief outlook of our statements for our algorithm, but see the detailed statements in later referred subsections.
@@
    (Theorem 2.4) There is a purely algebro-geometric $2$-step construction 222the following symbol $\rightsquigarrow$ means each step of the construction, and is not a map. of algebro-geometric version of minimum bubbles as affine varieties
@@
    Here, $\mathcal{X}^{\prime}_{\rm min,0},\mathcal{X}^{\prime\prime}_{\rm min,0}$ are both still $n$-dimensional affine log terminal varieties. The minimality morally means the mildness of singularity measured by the local normalized volumes (at the base points) as we clarify later. It is also parallel to minimality of the diverging order of the rescaling constants $\{c_{t}\}_{t}$ (cf., also [Sun23] and Definition 2.1).
    We call $\mathcal{X}^{\prime}_{\rm min,0}$ (resp., $\mathcal{X}^{\prime\prime}_{\rm min,0}$) (algebraic) semistable minimal bubble (resp., (algebraic) polystable minimal bubble) for simplicity. 333[Sun23] (resp., [dBS24]) conjectured such construction and called the semistable minimal bubble $\mathcal{X}^{\prime}_{\rm min,0}$ under different names as algebraic minimal bubble (resp., (minimal) weighted bubble). Note that our construction of $\mathcal{X}^{\prime\prime}_{\rm min,0}$ is also algebraic.
@@
    (Theorem 2.4, Theorem 2.11, Lemma 2.25) The first step construction
@@
We refer to these obtained (non-canonical) total spaces of the new degenerations to $\mathcal{X}^{\prime}_{\rm min,0}$ and $\mathcal{X}^{\prime\prime}_{\rm min,0}$ respectively as $\mathcal{X}^{\prime}_{\rm min}$ and $\mathcal{X}^{\prime\prime}_{\rm min}$.
By replacing $\mathcal{X}$ by $\mathcal{X}^{\prime\prime}_{\rm min}$ and do iteration of the above procedure ($\mathrm{i}$), we obtain deeper modifications of $\mathcal{X}$.
This whole process can be seen as a certain algorithm of improving the degenerations with some differential geometric motivation. Its double iterative nature may make look too complicated or even artificial, but it is unavoidable and indeed parallels the actual bubbling structure and hence we believe it is the nature of the arts in general. Nevertheless, it becomes simpler in many classical examples. For instance, in the $1$ or $2$-dimensional classical cases considered in [dBS24, §2.3, 3.3, 4.2], the second step of ($\mathrm{i}$) each time is trivial so that $\mathcal{X}^{\prime}_{\rm min,0}=\mathcal{X}^{\prime\prime}_{\rm min,0}$ recovers the differential geometric construction i.e., a certain pointed Gromov-Hausdorff limit (at the underlying affine varieties level) and the above construction of $\mathcal{X}^{\prime}_{\rm min,0}$ is canonical. Thus by iterations, we obtain the whole set of possible differential geometric bubbles along $\mathcal{X}\supset\sigma(\Delta)$ with the local metrics given in loc.cit. See Theorem 2.13.
Theorem 2.6 gives a subtle technical improvement of the above statements, including the following points. Firstly, we allow degenerations along more general bases. As a related advantage, the $1$-st step construction of the total space $\mathcal{X}^{\prime}_{\rm min}$ can be achieved in a more canonical manner i.e., as a more canonical birational modification of the original family, once we pass to a possibly non-finite type base change i.e., a certain Novikov type ring spectrum. To make sense of it, we use [Od24b] in which some algebraic (partly stack-theoretic) generalities to treat degenerations in “irrational direction” with higher dimensional algebraic base is prepared. We also compare our construction with more analytic and differential geometric constructions in [Sun23].
Since our construction is fairly general and look abstract first, it is designed after various explicit calculations as experiments. We do not write all the calculations but review one of the simplest examples after [dBS24] here. In this case, we only need a single process rather than the $2$-step processes i.e., $\mathcal{X}^{\prime}_{\rm min}=\mathcal{X}^{\prime\prime}_{\rm min}$, and they are achieved only after a finite base change of $\Delta$.
Example 1.2 (cf., [BKN89, Ban90], [dBS24, §3.3] and Theorem 2.13).
@@
We can consider a family of hyperKähler metrics on each $t$ by the Gibbons-Hawking ansatz with the base $\mathbb{R}^{3}$ and monopole points $x_{1},x_{2},x_{3},x_{4}$ and harmonic functions of the form $\frac{1}{2}\sum_{i}\frac{1}{|x-x_{i}|}$. In this case, [dBS24, §3.3] describes how it bubbles at $t\to 0$ around the origin, and coincides with the algebraic procedure Theorem 1.1, and is explicit as follows. (Indeed, our general construction is originally obtained as a generalization of their examples).
@@
after replacing $\frac{x}{t^{2}}$ as $x$, $\frac{y}{t^{2}}$ as $y$, $\frac{z}{t}$ as $z$ while keeping $t$. This is a degeneration to $xy=z^{3}(z-1)$ when $t\to 0$, with $A_{2}$-singularity at $(0,0,0,0)\in\Delta_{t}\times\mathbb{C}_{x,y,z}^{3}$. This admits a ALE hyperKähler orbifold metric, again obtained via the Gibbons-Hawking ansatz over $\mathbb{R}^{3}$ (cf., [dBS24, §3.3]).
@@
We make this kind of explicit study as Example 1.2 via coordinates more generally. That is, after the proof of Theorem 1.1 ($=$Theorems 2.4, 2.11), we propose more explicit ways of constructions of the candidates of bubbles, via coordinates/equations and valuations respectively in §2.4, §2.5. We conjecture that they coincide with the construction of Theorem 2.4. §2.2.5 and §2.3 may be of own independent interest, which are respectively the systematic treatment of negative weight deformation originated in [CH22, SJZ23] and an analogue of bubbling Calabi-Yau metric in the context of normalized Kähler-Ricci flow on Fano varieties. The latter would give a yet another approach to the properness of K-moduli of Fano varieties cf., Theorem 2.22. Further, in §3, we discuss more on the structure on the “bubble tree” formed by all possible bubbles.
@@ 2. Minimal bubble
2.1. Reviewing definitions
We work in Setup 1 in the introduction. In this section, we give algebraic approaches to the “minimum bubble” in the sense of [Sun23, dBS24], whereas the minimality roughly refers to that of the order of rescaling parameters $c_{t}$ for $t\to 0$. We review the definition for convenience.
@@ 2.1. Reviewing definitions
In the context of Setup 1 (a), its (differential geometric) minimal bubble refers to the pointed Gromov-Hausdorff limits of $(\sigma(t)\in\mathcal{X}_{t},c_{t}^{2}g_{t})$ for certain diverging sequence $c_{t}>0$, which is not a Riemannian metric cone, but whose tangent cone at infinity ([DS17, §3.4]) is the metric local tangent cone at $\sigma(0)\in X$, which we denote as $C_{\sigma(0)}(X)$.
Firstly, we discuss several a priori different algebraic approaches to it constructions, in different subsections, and will discuss the conjectural equivalence later. We crucially use the theory of local normalized volume of log terminal singularities due to [Li19, Liu18] (cf., also [Od15b]), with its origin in [DS17]. We briefly review the definition and properties we use here. Its smooth case can be traced back to much earlier [Mus02, dFEM04] (and also related [ELS03]) and more detailed review are available at [LLX20], [Od24a, §2] for instance.
@@
Then, the normalized volume $\widehat{{\rm vol}}_{x}(v)$ is defined as $A_{X}(v)^{n}{\rm vol}_{x}(v)$ by [Li19], where $A_{X}(v)$ is the log discrepancy in the generalized sense by [JM12]. Then the following is known by now. We do not make any substantial contribution in this paper but rather use in this paper. See more references in [LLX20], [Od24a, §2]. Morally speaking, this process is to modify a “locally K-stable” i.e., log terminal singularity (cf., [Od15b, GT23]) to cone, which is still (K-semi/poly)stable.
Theorem 2.3 (Scale up conification cf., [DS17, Li19, Liu18, Blu18, HS17, LX18, LWX21, LLX20, XZ21, Od24a]).
For any log terminal singularity $x\in X$, there is a unique quasi-monomial valuation $v=v_{X}$ with the center $x$ up to the $\mathbb{R}_{>0}$-action, which minimizes the normalized volume $\widehat{\operatorname{vol}}(-)$.
@@
    Then, the valuation $v_{X}$ coincides with the vanishing order function in [DS17, §3.2] for the local Kähler metric in the setup of loc.cit. We denote the rational rank of $v_{X}$ as $r$. Further, $\widehat{\rm vol}(v_{X})/n^{n}$ has a differential geometric meaning as the volume density of the link of the local metric tangent cone. We denote the minimized value $\widehat{\rm vol}(v_{X})$ as $\widehat{\rm vol}(x\in X)$ or $\widehat{\rm vol}_{X}(x)$, in this paper.
@@
    From more algebraic viewpoint, the graded ring
    ${\rm gr}_{v}(\mathcal{O}_{X,x}):=\oplus_{s\in{\rm Im}(v_{X})}(\{f\in\mathcal{O}_{X,x}\mid v(f)\geq s\}/\{f\in\mathcal{O}_{X,x}\mid v(f)>s\})$
    is a finite type $\mathbbm{k}$-algebra on which there is a natural action of the algebraic $\mathbbm{k}$-torus $T\simeq\mathbb{G}_{m}^{r}$ on a K-semistable Fano cone $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v}(\mathcal{O}_{X,x})$.
@@
We note that the last $\mathbb{Q}$-Gorensteinness assertion of the positive weight degeneration in ($\mathrm{iii}$) seems missing in literature but can be confirmed as follows. As in [Od24b, §2], take large enough homogeneous generating system $\overline{z}_{1},\cdots,\overline{z}_{l}$ of ${\rm gr}_{v}(\mathcal{O}_{X,x})$ with degrees $w_{1},\cdots,w_{l}$, and lift to $\Gamma(\mathcal{O}_{X})$ as $\overline{z}_{1},\cdots,\overline{z}_{l}$ which we suppose to give an embedding $W\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}$ (otherwise add more $\overline{z}_{i}$s). Then we take rational approximation $\xi^{\prime}$ of $\xi:=(w_{1},\cdots,w_{l})$ and corresponding positive weight test configuration as $W_{\xi^{\prime}}$ over $\mathbb{A}^{1}$. For any $\xi^{\prime}$ close enough to $\xi$, we see that the central fiber is again $W$ just with a perturbed vector field induced on it. Note that there is a natural relative compactification $W_{\xi^{\prime}}\subset\overline{W_{\xi^{\prime}}}\to\mathbb{A}^{1}$ by adding the quotient by $\mathbb{G}_{m}$ for $\xi^{\prime}$-direction, which is automatically a $\mathbb{Q}$-Gorenstein family by the arguments of [OSS16, (proof of) Lemma 2.4]. Since this holds for any $\xi^{\prime}$ close enough to $\xi$, the map from the toric base $U_{\tau}$ of the positive weight degeneration is in the $\mathbb{Q}$-Gorenstein deformation component so that we obtain the assertion.
2.2. Construction via deformation theory
We use theorems in [Od24b] after [Lan75, HL14, AHLH23] to algebraic construction of the conjectural (minimal) bubble and discuss it. We also discuss reformulation of negative weight deformations in the sense of [CH22], after [Od24b], in which we also called it scale down deformation.